### v1

@@ Abstract
Hybrid models combine mechanistic ODE-based dynamics with flexible and expressive neural network components. Such models have grown rapidly in popularity, especially in scientific domains where such ODE-based modeling offers important interpretability and validated causal grounding (e.g., for counterfactual reasoning). The incorporation of mechanistic models also provides inductive bias in standard blackbox modeling approaches, critical when learning from small datasets or partially observed, complex systems. Unfortunately, as hybrid models become more flexible, the causal grounding provided by the mechanistic model can quickly be lost. We address this problem by leveraging another common source of domain knowledge: ranking of treatment effects for a set of interventions, even if the precise treatment effect is unknown. We encode this information in a causal loss that we combine with the standard predictive loss to arrive at a hybrid loss that biases our learning towards causally valid hybrid models. We demonstrate our ability to achieve a win-win — state-of-the-art predictive performance and causal validity — in the challenging task of modeling glucose dynamics during exercise.
@@ 1 Introduction
Blackbox sequence models have demonstrated extraordinary performance in numerous fields, including natural language processing, robotics, and demand forecasting. Popular methods include recurrent neural networks (RNNs), such as long short-term memory networks (LSTMs) (Hochreiter and Schmidhuber, 1997) and gated recurrent units (GRUs) (Chung et al., 2014); (temporal) convolutional neural networks (CNNs) (Lea et al., 2016; Bai et al., 2018; Shi et al., 2023); Transformer models, including Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021) variants; and state space sequence models such as S4 (Gu et al., 2021) and its diagonal approximation S4D (Gu et al., 2022). Despite their transformative role in many fields, the application of these models to various scientific domains has encountered challenges. An obvious hurdle is the limited size of some scientific datasets. But even in the presence of lots of data, these methods are still crippled by the incomplete nature of what can be measured. This problem is exacerbated by the fact that most data is observational: blackbox models are adept at identifying associations leading to good predictions, but may not learn causally coherent models. For example, in our T1D setting, insulin delivery frequently coincides with a planned meal or snack, which initially causes glucose to rise; so the model may incorrectly infer that insulin causes glucose to rise, when the reality is the opposite.
@@
Sophisticated mechanistic models—specified via a set of ordinary differential equations (ODEs)—remain a preferred method in many scientific domains, as they capture lab-validated or otherwise known physical or physiological causal properties of the system. Examples include cardiac and renal modeling (Hilgemann and Noble, 1987; ten Tusscher et al., 2004; Marsh et al., 2005), immunology and viral kinetics (Perelson et al., 1996; Canini and Perelson, 2014; Passini et al., 2017), epidemiology (He et al., 2020), neural circuits (Hodgkin and Huxley, 1952; Ladenbauer et al., 2019), and pharmacokinetics (Holz and Fahr, 2001). The UVA/Padova simulator (Man et al., 2014) of insulin-glucose dynamics we consider in this paper has been FDA approved as a substitute for pre-clinical trials in the development of artificial pancreas algorithms. Of note, however, is that sophisticated mechanistic models tend to add states to capture more intricate, non-Markovian dynamics such as delays, while maintaining the tractability of linear ODEs, leading to over-parameterization, and as a result, nonidentifiability. The latest version of the UVA/Padova simulator has over 30 states. While highly parameterized, the solution space of these models fails to capture important dynamics observed in real-world data; moreover, their parameter inference is highly sensitive. Further limiting this class of models is the typical assumption of a fixed mechanistic structure with a static set of simulator parameters. In T1D, many unobserved or partially-observed time-varying factors affect glycemic responses, including stress, hormone cycles, sleep, and activity levels (cf., Wellen et al., 2005). Likewise, the causal mechanism governing glycemic responses can vary—e.g., during certain types of physical activity, multiple underlying processes (not modeled in UVA/Padova) are activated based on the available energy sources (McArdle et al., 2006).
@@
Given these limitations, we focus on an alternative approach that hybridizes machine learning (ML) with domain knowledge encoded in mechanistic models. These hybrid models have gained traction across the natural and physical sciences (Willard et al., 2022), while being given different names and interpretations such as graybox modeling (Rico-Martinez et al., 1994), physics-informed machine learning (Karniadakis et al., 2021), universal differential equations (Rackauckas et al., 2020) and neural closure learning (Gupta and Lermusiaux, 2021). The core idea of hybrid modeling is to infuse domain inductive bias such that the trained model can gain the best of both worlds–mechanistic rigor with the flexibility and expressivity of deep learning. The hope is that hybrid methods not only allow one to solve complex modeling problems with improved precision and accuracy (Pathak et al., 2018; Willard et al., 2022), but also reduce the demand on data (Rackauckas et al., 2020), enhance reliability and robustness (Didona et al., 2015), and make the ML algorithms interpretable (Karniadakis et al., 2021; Du et al., 2019). Examples of successes of hybrid modeling in healthcare can be found in Qian et al. (2021); Sottile et al. (2021); Hussain et al. (2021).
@@
Indeed, in Fig. 1, we see the important inductive bias such mechanistic models provide. On the one hand, the pure mechanistic model makes poor predictions due to its brittle nature while on the other hand, the flexible blackbox models are subject to overfitting. The hybrid models outperform either alternative in terms of prediction error.
@@
Unfortunately, as we also see in Fig. 1, hybrid models can quickly lose their valid causal grounding as more and more flexible modeling components are deployed. When tasked with selecting the intervention with the largest treatment effect amongst a set of three counterfactual simulations (see Sec. 5), classification error for hybrid models gradually deteriorates as the model becomes more blackbox. Amongst the blackbox models themselves, we see that the majority of black-box models have a classification error rate above $67\%$, the error rate of random guessing.
[…]
@@
Figure 1: Prediction RMSE and its standard error (left) vs. 10/50/90-th (lower/bar/upper) percentile classification error rate (right) for (1) UVA/Padova mechanistic model (UVA); (2) hybrid models: latent parameter dynamics (LP), latent parameter dynamics + state closure (LPSC), MNODE; and (3) blackbox models: neural ODE (BNODE), temporal convolutional network (TCN), LSTM, Transformer (Trans) and S4D, all trained on pure predictive loss ($\alpha=0$). The dashed gray line in the second figure corresponds to the causal classification error rate of random guessing ($2/3$).
@@ 2 The Hybrid Modeling Spectrum
Figure 2: Visualization of hybrid modeling spectrum for a simple mechanistic model of insulin-glucose dynamics. From left to right: (1) mechanistic model; (2) latent parameter dynamics; (3) MNODE; (4) neural ODE. Dependencies between states maintained until (4), with state dynamics increasingly flexible.
@@ 2.1 Mechanistic Models
On the far left hand side of our hybrid modeling spectrum in Fig. 2 is the pure mechanistic model specified via a set of ordinary differential equations (ODE),
@@
where $s$ is a vector representing the state, and $\beta$ is a vector representing the simulator parameters. In applications where interventions or other external controls represented by a vector $x$ are being applied to the observed dynamical system, we obtain a controlled ordinary differential equation (CDE),
@@ 2.2 Neural ODE
In the absence of mechanistic domain knowledge, neural ODEs (Rico-Martinez et al., 1992; Weinan, 2017; Haber and Ruthotto, 2017; Chen et al., 2018; Kidger, 2022) have been proposed as a blackbox approach to learning such a dynamical system:
@@
Here, $f$ denotes a neural network with parameters $\theta$. Neural ODEs represent the far right side of the hybrid modeling spectrum in Fig. 2. These methods have proven useful in many dynamical system modeling and simulation tasks, especially in the sciences where ODEs are a standard language for describing systems (Qian et al., 2021; Lu et al., 2021; Owoyele and Pal, 2022; Asikis et al., 2022; Li et al., 2022).
@@ 2.3 Hybrid Models
Levine and Stuart (2022) and Karniadakis et al. (2021) provide related frameworks for defining the space of hybrid models. The above schemes, and those further described in Levine and Stuart (2022); Karniadakis et al. (2021), are not mutually exclusive and are often combined (Qian et al., 2021; Wu et al., 2022; Wang et al., 2022).
@@
We address this challenge in Sec. 3.2 by introducing a hybrid loss that mixes predictive performance with causal validity. For the latter, we again lean on domain knowledge (as we did in our use of a mechanistic model), but this time cast as knowledge about the direction of treatment effects. In other words, we presume that we know in advance that applying, e.g., $2x$ instead of $x$ to the system will increase (or decrease) the value of a target evaluation metric that depends on the observed dynamic (counterfactual) response. Encoding this information in the loss function helps us achieve a win-win of increased modeling flexibility while biasing the learning towards solutions that match our causal knowledge.
@@ 3.1 Limitations of Hybrid Models
To illustrate the potential for hybrid models to lose causal validity, take the MNODE model of Eq. (5) where $c_{1}=0$. Relying solely on the dependency graph between states while learning state dynamics can fail to distinguish parameters leading to correct directions of treatment effects. Consider data generated by a mechanistic ODE $ds/dt=-as+bx+c$, where $a,b,c$ are positive. Here the rate of change of $s$ is monotonically increasing with respect to $x$ and therefore an intervention that increases $x$ should have a positive treatment effect on $s$. However, after turning these dynamics into MNODE, we instead have $ds/dt=f(s,x;\theta);$ here, we lose the monotonicity with respect to $x$, unless a monotone neural network is intentionally used—which would severely constrain the flexibility of MNODE, the original motivation for adopting this modeling approach.
One might argue that while not explicitly encoded in the architecture, neural networks should be able to learn relatively simple signals like monotonicity from the data. Unfortunately, this is not the case when learning from noisy, observational datasets. For example, in T1D, the dosing of bolus insulin (negative effect on glucose) frequently occurs in close succession with a planned meal (positive effect). Due to the confounding effect of insulin, the hybrid model may incorrectly infer that if an individual consumes carbohydrates, glucose drops; see Fig. 3. The resulting dilemma is, while we do not want to make simplifying assumptions like linearity and monotonicity, we want to encourage our hybrid models to learn dynamics and interactions consistent with known signs of treatment effects.
@@ 3.2 The Hybrid Objective
Let $X(\mathcal{T})=\left\{x(t);t\in\mathcal{T}\right\}$ be the set of control inputs at times in $\mathcal{T}$, $\mathcal{T}\subset\mathbb{R}$. In our diabetes example, $x$ can represent interventions such as carbohydrate intake or insulin dosing for a single patient. In our preceding development, we describe models for evolution of the full state $[s(t);z(t)]$. Here we assume we have a partial observation of that state; for example, in our diabetes example, we observe glucose measurements via CGM. We denote this partial state observation by $y(t)$, and we assume without loss of generality that $y(t)=H[s(t);z(t)]$ for some projection matrix $H$ (which can be the identity matrix, $I$, if we have complete observations). We then define $Y(\mathcal{T})=\left\{y(t);t\in\mathcal{T}\right\}$ to be the corresponding set of (partial) observations over $\mathcal{T}$.
We assume we have some context window $\mathcal{T}_{\mathrm{context}}$ of controls and observations, $(X(\mathcal{T}_{\mathrm{context}}),\allowbreak Y(\mathcal{T}_{\mathrm{context}}))$. We are interested in the behavior of $y$ over a prediction window $\mathcal{T}_{\mathrm{pred}}$, represented by $Y(\mathcal{T}_{\mathrm{pred}})$, given future controls $X(\mathcal{T}_{\mathrm{pred}})$. A model $M$ takes as input $(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}})$, $X(\mathcal{T}_{\mathrm{pred}}))$ and produces an output trajectory of $y$ over the prediction window. Below we use $M^{\dagger}$ to represent the unknown ground truth model; this is the model that generated the observations $Y(\mathcal{T}_{\mathrm{pred}})$. We use $\hat{M}$ to denote a fitted model. Throughout we use $\hat{Y}(\mathcal{T}_{\mathrm{pred}})=\left\{\hat{y}(t);t\in\mathcal{T}_{\mathrm{pred}}\right\}$ to denote our predicted output from such a fitted model.
@@
In many settings, we have a value function of our observed trajectory $u(Y(\mathcal{T}_{\mathrm{pred}}))$, that computes a meaningful scalar-valued output from a sequence input. For example, in diabetes, this value might correspond to the average glucose level over $\mathcal{T}_{\mathrm{pred}}$. For fixed context $X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}})$, we consider a range of interventions $X^{(i)}(\mathcal{T}_{\mathrm{pred}})$, $i=1,\ldots,K$, that differ from the observed $X(\mathcal{T}_{\mathrm{pred}})$. The ground truth model would produce observations $Y^{(i)}(\mathcal{T}_{\mathrm{pred}})=M^{\dagger}(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}),X^{(i)}(\mathcal{T}_{\mathrm{pred}}))$ for the $i$’th intervention. If we could observe these counterfactuals, then we could compute the causal effect of each intervention:
@@
Domain knowledge provides valuable qualitative guidance on the direction and ordering of these treatment effects: for example, we know that holding other context constant, increasing insulin dosing should lower glucose levels, and more insulin leads to greater reductions. Unfortunately, often these causal effects cannot be precisely quantitatively estimated. For example, in managing diabetes, exact estimation of glucose levels is exceedingly challenging even in highly controlled settings.
Instead, in our causal loss, we evaluate whether the model can identify the intervention with the maximum value, rather than precise estimation of causal effects. We define:
*$\mathcal{I}^{*}=\text{argmax}_{i}\tau^{(i)}=\text{argmax}_{i}u(Y^{(i)}(\mathcal{T}_{\mathrm{pred}})).$
In principle, we can compare this intervention to an estimated maximum value intervention under the trained model $\hat{M}$, i.e.,
@@
where $\hat{Y}^{(i)}(\mathcal{T}_{\mathrm{pred}})=\hat{M}(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}),$ $X^{(i)}(\mathcal{T}_{\mathrm{pred}}))$ for the $i$’th intervention.
Of course, to enable model training, we need a loss function that admits a gradient. Accordingly, rather than computing the exact maximum value intervention under the estimated model, we compute the softmax vector:
@@
For a given $\alpha\in[0,1]$, we define our hybrid loss as:
@@
where we have overloaded notation to interpret $L_{\mathrm{pred}}$ and $L_{\mathrm{causal}}$ as the average predictive and causal losses respectively over all sequences and scenarios. Note that when $\alpha=0$, model fitting focuses entirely on prediction; when $\alpha=1$, model fitting focuses entirely on causal validity.
[…]
A critical question in fitting hybrid ODE models is the initial condition $(s_{0}=s(t_{0}),z_{0}=z(t_{0}))$, where $t_{0}$ is the beginning of our prediction window. Two primary approaches exist for performing initial state estimation jointly with hybrid ODE parameter learning: (1) statistical state estimation methods (Chen et al., 2022; Brajard et al., 2021; Ribera et al., 2022; Levine and Stuart, 2022) and (2) blackbox encoder models (Chen et al., 2018). We take the latter approach, viewing initial state estimation through the lens of sequence-to-sequence (seq2seq) modeling.
In particular, we produce the predicted sequence $\hat{Y}(\mathcal{T}_{\mathrm{pred}})$ using $X(\mathcal{T}_{\mathrm{pred}})$ and the initial condition $(s_{0},z_{0})$ encoded from the context data $(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}))$. We consider a general blackbox encoder for the initial conditions:
$(s_{0},z_{0})=M_{\mathrm{context}}\big{(}X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}});\theta_{\mathrm{context}}\big{)},$ (9)
where $\theta_{\mathrm{context}}$ are the parameters of the encoder.
@@
$\hat{y}\big{(}t_{k}\big{)}=H\cdot\mbox{Integrate}(M_{\mathrm{pred}},\mathcal{T}_{\mathrm{pred}};X(\mathcal{T}_{\mathrm{pred}}),s_{0},z_{0},\theta_{\mathrm{pred}}).$ (10)
[…]

### v2

@@ Abstract
Hybrid models composing mechanistic ODE-based dynamics with flexible and expressive neural network components have grown rapidly in popularity, especially in scientific domains where such ODE-based modeling offers important interpretability and validated causal grounding (e.g., for counterfactual reasoning). The incorporation of mechanistic models also provides inductive bias in standard blackbox modeling approaches, critical when learning from small datasets or partially observed, complex systems. Unfortunately, as the hybrid models become more flexible, the causal grounding provided by the mechanistic model can quickly be lost. We address this problem by leveraging another common source of domain knowledge: ranking of treatment effects for a set of interventions, even if the precise treatment effect is unknown. We encode this information in a causal loss that we combine with the standard predictive loss to arrive at a hybrid loss that biases our learning towards causally valid hybrid models. We demonstrate our ability to achieve a win-win, state-of-the-art predictive performance and causal validity, in the challenging task of modeling glucose dynamics post-exercise in individuals with type 1 diabetes.
Machine Learning, ICML
@@ 1 Introduction
Blackbox sequence models have demonstrated extraordinary performance in numerous fields, ranging from natural language processing to forecasting. Popular methods include recurrent neural networks (RNNs), such as long short-term memory networks (LSTMs) (Hochreiter & Schmidhuber, 1997) and gated recurrent units (GRUs) (Chung et al., 2014); (temporal) convolutional neural networks (CNNs) (Lea et al., 2016; Bai et al., 2018; Shi et al., 2023); Transformer models, including Informer (Zhou et al., 2021) and Autoformer (Wu et al., 2021); and state space sequence models, such as S4 (Gu et al., 2022b). Despite their transformative role in many fields, the application of these models to various scientific domains has encountered challenges. An obvious hurdle is the limited size of some scientific datasets. But even in the presence of lots of data, these methods are still crippled by the incomplete nature of what can be measured. This problem is exacerbated by the fact that most data is observational: blackbox models are adept at identifying associations leading to good predictions, but may not learn causally coherent models. For example, in our T1D setting, insulin delivery frequently coincides with a planned meal or snack, which initially causes glucose to rise; the model may incorrectly infer that insulin causes glucose to rise, when the reality is the opposite.
@@
Sophisticated mechanistic models—specified via a set of ordinary differential equations (ODEs)—remain a preferred method in many scientific domains, as they capture lab-validated or otherwise known physical or physiological causal properties of the system. Examples include cardiac and renal modeling (Hilgemann & Noble, 1987; ten Tusscher et al., 2004; Marsh et al., 2005; Passini et al., 2017), immunology and viral kinetics (Perelson et al., 1996; Canini & Perelson, 2014), epidemiology (He et al., 2020), neural circuits (Hodgkin & Huxley, 1952; Ladenbauer et al., 2019), and pharmacokinetics (Holz & Fahr, 2001). The UVA/Padova simulator (Man et al., 2014) of insulin-glucose dynamics we consider in this paper has been FDA approved as a substitute for pre-clinical trials in the development of artificial pancreas algorithms. Notably, sophisticated mechanistic models often introduce additional states to capture intricate, non-Markovian dynamics such as delays. While this approach preserves the tractability of linear ODEs, it can lead to over-parameterization. The latest version of the UVA/Padova simulator has over 30 states. Yet, these models fail to capture important dynamics observed in real-world data; moreover, their parameter inference is highly sensitive. Further limiting this class of models is the typical assumption of a fixed mechanistic structure with a static set of simulator parameters. In T1D, many unobserved or partially-observed time-varying factors affect glycemic responses, including stress, hormone cycles, sleep, and activity levels (cf., Wellen et al., 2005). Likewise, the causal mechanism governing glycemic responses can vary—e.g., during certain types of physical activity, multiple underlying processes (not modeled in UVA/Padova) are activated based on available energy sources (McArdle et al., 2006).
@@
Given these limitations, we focus on an alternative approach that hybridizes machine learning (ML) with domain knowledge encoded in mechanistic models. These hybrid models have gained traction across the natural and physical sciences (Willard et al., 2022), while being given different names and interpretations such as graybox modeling (Rico-Martinez et al., 1994), physics-informed machine learning (Karniadakis et al., 2021), universal differential equations (Rackauckas et al., 2020) and neural closure learning (Gupta & Lermusiaux, 2021). The core idea of hybrid modeling is to infuse domain inductive bias such that the trained model can gain the best of both worlds—mechanistic rigor with the flexibility and expressivity of deep learning. The hope is that hybrid methods not only allow one to solve complex modeling problems with improved precision and accuracy (Pathak et al., 2018; Willard et al., 2022), but also reduce the demand on data (Rackauckas et al., 2020), enhance reliability and robustness (Didona et al., 2015), and make the ML algorithms interpretable (Karniadakis et al., 2021; Du et al., 2019). Examples of successes of hybrid modeling in healthcare can be found in Qian et al. (2021); Sottile et al. (2021); Hussain et al. (2021).
@@
Indeed, in Fig. 1 (left), we see the important inductive bias such mechanistic models provide. On the one hand, the pure mechanistic model makes poor predictions due to its brittle nature while on the other hand, the flexible blackbox models are subject to overfitting. The hybrid models generally outperform either alternative in terms of prediction error.
@@
Unfortunately, as we also see in Fig. 1 (right), hybrid models can quickly lose their valid causal grounding as more and more flexible modeling components are deployed. When tasked with selecting the intervention with the largest treatment effect amongst a set of three counterfactual simulations (see Sec. 5), classification error for hybrid models gradually deteriorates as the model becomes more blackbox. Amongst the hybrid and blackbox models, we see that the majority have a classification error rate statistically indistinguishable from random guessing, 67%.
To address these challenges, we leverage another common source of domain knowledge: rankings of various treatment effects. While domain knowledge is often insufficient to specify expected treatment effects under counterfactually-applied interventions, we are still often able to make quantitative claims about relative magnitudes of different treatment effects. For example, while we may not know the precise treatment effect of eating a small salad versus a whole birthday cake (all else held constant), we do know the sign and relative scale of treatment effect. As this prior knowledge is challenging to encode in the hybrid model itself, we introduce a causal loss function that encourages our learned model to perform well in these comparison tasks. Our hybrid loss is a convex combination of predictive loss and causal loss. When applied to training hybrid models, we refer to the overall method as hybrid2 neural ODE causal modeling (H2NCM). In a real-world task of guiding individuals with T1D on the impact of interventions so they can safely exercise, we show that our H2NCM approach achieves the best of both worlds: state-of-the-art predictive accuracy with causal validity.
@@
Figure 1: Prediction RMSE and standard error (left) vs. 10/50/90-th (lower/bar/upper) percentile classification error rate (right) for (1) UVA/Padova mechanistic model (UVA); (2) hybrid models: latent parameter dynamics (LP), latent parameter dynamics + state closure (LPSC), MNODE; and (3) blackbox models: neural ODE (BNODE), temporal convolutional network (TCN), LSTM, Transformer (Trans) and S4D, all trained on pure predictive loss ($\alpha=0$). The dashed gray line in the second figure corresponds to the causal classification error rate of random guessing ($2/3$).
@@ 2 The Hybrid Modeling Spectrum
Figure 2: Visualization of hybrid modeling spectrum for a simple mechanistic model of insulin-glucose dynamics. From left to right: (1) mechanistic model; (2) latent parameter dynamics; (3) MNODE; (4) neural ODE. Dependencies between states are maintained until (4), with state dynamics increasingly flexible.
@@ 2.1 Mechanistic Models
On the far left-hand side of our hybrid modeling spectrum in Fig. 2 is the pure mechanistic model specified via a set of ordinary differential equations (ODE),
@@
where $s$ is a vector representing the state, and $\beta$ is a vector representing the simulator parameters. In applications where interventions or other external controls, $x$, are being applied to the observed dynamical system, we obtain a controlled ordinary differential equation (CDE),
@@ 2.2 Neural ODE
In the absence of mechanistic domain knowledge, neural ODEs (Rico-Martinez et al., 1992; Weinan, 2017; Haber & Ruthotto, 2017; Chen et al., 2018; Kidger, 2021) have been proposed as a blackbox approach to learning such a dynamical system:
@@
Here, $f$ denotes a neural network with parameters $\theta$. Neural ODEs represent the far right side of the spectrum in Fig. 2. These methods have proven useful in many dynamical system modeling and simulation tasks, especially in the sciences where ODEs are a standard language for describing systems (Qian et al., 2021; Lu et al., 2021; Owoyele & Pal, 2022; Asikis et al., 2022; Li et al., 2022).
@@ 2.3 Hybrid Models
Yin et al. (2021), Levine & Stuart (2022), and Karniadakis et al. (2021) provide related frameworks for defining the space of hybrid models. The above schemes, and those further described in Levine & Stuart (2022); Karniadakis et al. (2021); Yin et al. (2021), are not mutually exclusive and are often combined (Qian et al., 2021; Wu et al., 2022; Wang et al., 2022).
@@ 3 Hybrid2 Modeling
We address this challenge in Sec. 3.2 by introducing a hybrid loss that mixes predictive performance with causal validity. For the latter, we again lean on domain knowledge (as we did in our use of a mechanistic model), but this time cast as knowledge about the direction of treatment effects. In other words, we presume that we know in advance that applying, e.g., $2x$ instead of $x$ to the system will increase (or decrease) the value of a score function that depends on the observed dynamic (counterfactual) response. Encoding this information in the loss function helps us achieve a win-win of increased modeling flexibility while biasing the learning towards solutions that match our causal knowledge.
@@ 3.1 Limitations of Hybrid Models
To illustrate the potential for hybrid models to lose causal validity, take the MNODE model of Eq. (5) where $c_{1}=0$. Relying solely on the mechanistic dependency graph between states while learning state dynamics can fail to distinguish parameters leading to correct directions of treatment effects. Consider data generated by a mechanistic ODE $ds/dt=-as+bx+c$, where $a,b,c$ are positive. Here, the rate of change of $s$ is monotonically increasing with $x$ and therefore an intervention that increases $x$ should have a positive treatment effect on $s$. However, after turning these dynamics into MNODE, we instead have $ds/dt=f(s,x;\theta);$ here, we lose the monotonicity with respect to $x$, unless a monotone neural network is intentionally used—which would severely constrain the flexibility of MNODE, the original motivation for adopting this modeling approach.
One might argue that while not explicitly encoded in the architecture, neural networks should be able to learn relatively simple signals like monotonicity from the data. Unfortunately, this is not the case when learning from noisy, observational datasets. For example, in T1D, the dosing of bolus insulin (negative effect on glucose) frequently occurs in close succession with a planned intake of carbohydrates (positive effect). Due to the confounding effect of insulin, the hybrid model may incorrectly infer that if an individual consumes carbohydrates, glucose drops; see Fig. 3. We provide an exploration of this effect in Sec. 5.4. The resulting dilemma is, while we do not want to make simplifying assumptions like linearity and monotonicity, we want to encourage our hybrid models to learn dynamics and interactions consistent with known signs of treatment effects.
@@ 3.2 The Hybrid Objective
Let $X(\mathcal{T})=\left\{x(t);t\in\mathcal{T}\right\}$ be the set of control inputs at times in $\mathcal{T}$, $\mathcal{T}\subset\mathbb{R}$. In our diabetes example, $x$ can represent interventions such as carbohydrate intake or insulin dosing for a single patient. In our preceding development, we described models for the evolution of the full state $[s(t);z(t)]$. Here we assume we have a partial observation of that state; for example, in our diabetes case study, we observe glucose measurements via a continuous glucose monitor (CGM). We denote this partial state observation by $y(t)$ and assume without loss of generality that $y(t)=H[s(t);z(t)]$ for some projection matrix $H$ (which can be the identity matrix, $I$, if we have complete observations). We then define $Y(\mathcal{T})=\left\{y(t);t\in\mathcal{T}\right\}$ as the corresponding set of (partial) observations over $\mathcal{T}$.
We assume a context window $\mathcal{T}_{\mathrm{context}}$ of controls and observations, $(X(\mathcal{T}_{\mathrm{context}}),\allowbreak Y(\mathcal{T}_{\mathrm{context}}))$. We are interested in the behavior of $y(t)$ over a prediction window $\mathcal{T}_{\mathrm{pred}}$, represented by $Y(\mathcal{T}_{\mathrm{pred}})$, given future controls $X(\mathcal{T}_{\mathrm{pred}})$. A model $M$ takes as input $(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}})$, $X(\mathcal{T}_{\mathrm{pred}}))$ and produces an output trajectory of $y$ over the prediction window. Below we use $M^{\dagger}$ to represent the unknown ground truth model; this is the model that generated the observations $Y(\mathcal{T}_{\mathrm{pred}})$. We use $\hat{M}$ to denote a fitted model. Throughout we use $\hat{Y}(\mathcal{T}_{\mathrm{pred}})=\left\{\hat{y}(t);t\in\mathcal{T}_{\mathrm{pred}}\right\}$ to denote our predicted output from such a fitted model.
@@
To incorporate causal validity into the loss function, we develop a causal loss that examines counterfactual simulations rather than predictions. For fixed context $X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}})$, we consider a range of hypothetical interventions $X^{(i)}(\mathcal{T}_{\mathrm{pred}})$, $i=1,\ldots,K$, that differ from the observed $X(\mathcal{T}_{\mathrm{pred}})$. The ground truth model would produce observations $Y^{(i)}(\mathcal{T}_{\mathrm{pred}})=M^{\dagger}(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}),X^{(i)}(\mathcal{T}_{\mathrm{pred}}))$ for the $i$’th intervention. If we could observe these counterfactuals, then we could compute the causal effect of each intervention:
@@
Here, $u$ is a score function that computes a meaningful scalar-valued output from a sequence input. For example, in diabetes, this score might correspond to the average glucose level over $\mathcal{T}_{\mathrm{pred}}$. In other settings, the relevant quantity may be related to extremes, in which case $u$ may compute, e.g., the maximum or minimum value of the sequence.
In many settings with complex dynamics or only very partial observations, such as in modeling glycemic response, the latent sources of variation (e.g., sleep, stress, physiology, etc.) render attempts to precisely estimate treatment effects nearly impossible. Importantly, in our setting the clinically relevant quantity for guiding patient care is the ordering of interventions; the same would be true in a multitude of settings where the model is guiding decision making. This ordinal domain knowledge is often readily available; for example, we know that holding other context constant, increasing insulin dosing should lower glucose levels, and increasingly larger insulin doses lead to greater reductions in glucose levels. Even such ordinal causal grounding is absent in purely predictive models.
To bias our training towards models that capture this ordinal causal domain knowledge, we introduce a causal loss where we evaluate whether the model can indentify the intervention with the maximum score. We define:
$\mathcal{I}^{*}=\text{argmax}_{i}\tau^{(i)}=\text{argmax}_{i}u(Y^{(i)}(\mathcal{T}_{\mathrm{pred}})).$
In principle, we can compare this intervention to an estimated maximum score intervention under the trained model $\hat{M}$, i.e.,
@@
where for the $i$’th intervention we have $\hat{Y}^{(i)}(\mathcal{T}_{\mathrm{pred}})=\hat{M}(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}),$ $X^{(i)}(\mathcal{T}_{\mathrm{pred}}))$.
Of course, to enable model training, we need a loss function that admits a gradient. Accordingly, rather than computing the exact maximum score intervention under the estimated model, we compute the softmax vector:
@@
For $\alpha\in[0,1]$, we define our hybrid loss as:
@@
where we have overloaded notation to interpret $L_{\mathrm{pred}}$ and $L_{\mathrm{causal}}$ as the average predictive and causal losses, respectively, over all sequences and scenarios. Note that when $\alpha=0$, model fitting focuses entirely on prediction; when $\alpha=1$, model fitting focuses entirely on causal validity.
Model Fitting: The H2NCM Approach
A critical question in fitting hybrid ODE models is the initial condition $(s_{0}=s(t_{0}),z_{0}=z(t_{0}))$, where $t_{0}$ is the beginning of our prediction window. Two primary approaches exist for performing initial state estimation jointly with hybrid ODE parameter learning: (1) statistical state estimation methods (Chen et al., 2022; Brajard et al., 2021; Ribera et al., 2022; Levine & Stuart, 2022) and (2) blackbox encoder models (Chen et al., 2018). We take the latter approach, viewing initial state estimation through the lens of sequence-to-sequence (seq2seq) modeling.
In particular, we produce the predicted sequence $\hat{Y}(\mathcal{T}_{\mathrm{pred}})$ using $X(\mathcal{T}_{\mathrm{pred}})$ and the initial condition $(s_{0},z_{0})$ encoded from the context data $(X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}}))$. We consider a general blackbox encoder $M_{\mathrm{context}}$ for the initial conditions:
$(s_{0},z_{0})=M_{\mathrm{context}}\big{(}X(\mathcal{T}_{\mathrm{context}}),Y(\mathcal{T}_{\mathrm{context}});\theta_{\mathrm{context}}\big{)},$
where $\theta_{\mathrm{context}}$ are the parameters of $M_{\mathrm{context}}$.
@@
$\hat{y}\big{(}t_{k}\big{)}=H\cdot\mbox{Integrate}(M_{\mathrm{pred}},\mathcal{T}_{\mathrm{pred}};X(\mathcal{T}_{\mathrm{pred}}),s_{0},z_{0},\theta_{\mathrm{pred}}).$
Here, $H$ is an indicator matrix selecting the observed states. The parameters $\theta_{\mathrm{context}}$ and $\theta_{\mathrm{pred}}$ are jointly optimized to obtain a fitted model $\hat{M}$ that minimizes the hybrid loss of Eq. (8). We refer to this approach of training such a model by optimization of hybrid loss as H2NCM. [^2] Our implementation of H2NCM is available at https://github.com/bobjz/H2NCM.